A firm has a market value of equity of $30,000 . It borrows $7500 at a cost of 8%. If the firm’s assets have a cost of capital of 15%, what is the firmʹs cost of equity capital? Assume no taxes.

6.70%

20.10%

23.45%

16.75%
Answer #1

Answer is 16.75%

WACC = Weight of debt * Cost of debt + Weight of equity * Cost of equity

Weight of Debt = 7500/(30000 + 7500) = 20%

Weight of Equity = 30000/(30000 + 7500) = 80%

15% = 20% * 8% + 80% * Cost of Equity

15% = 1.6% + 80% * Cost of Equity

13.4% = 80% * Cost of Equity

Cost of Equity = 13.4%/80% = 16.75%